OpenStudy (anonymous):

Find the length of the base of a square pyramid if the volume is 576 cubic inches and has a height of 3 inches. 4 inches 8 inches 16 inches 24 inches

OpenStudy (jack1):

Volume of a pyramid = ( Length x Width x Height ) / 3 if the base is a square: Length = width so Volume of a pyramid = ( Length x Length x Height ) / 3 \(\Large V = \frac{L \times L \times H}3\) if you have the volume and the height... can you solve this for Length?

OpenStudy (anonymous):

would the length be 576 and the height 3?

OpenStudy (jack1):

no, V = 576 and h = 3

OpenStudy (anonymous):

oh ok sorry um.. im just wondering idk what length would be because i only have volume and height

OpenStudy (jack1):

sók... yep, we're trying to solve the length if u plug the values for volume and height into the above equation... you should be left with L^2 = ... a number

OpenStudy (anonymous):

ok 576/3 would be 192 right ??

OpenStudy (jack1):

\(\Large V = \frac{L \times L \times H}3\) \(\Large 576 = \frac{L \times L \times 3}3\) \(\Large 576 = \frac{L \times L \times \color{red}{3}} {\color{red}{3}}\) \(\Large 576 = L \times L \) \(\Large 576 = L^2 \) \(\Large \sqrt{576} = \sqrt{L^2} \) \(\Large \sqrt{576} = L \)

OpenStudy (jack1):

sorry does this make it clearer tho?

OpenStudy (anonymous):

a little bit im still kinda confused bu ti some what understand

OpenStudy (jack1):

does the picture help? where abouts are you getting stuck and I'll try to explain it better

OpenStudy (anonymous):

i was just confused at the part of where we had to find the length. but that chart u showed does help a little

OpenStudy (jack1):

cools... so how'd u go at solving, did u get an answer for L = ...?

OpenStudy (anonymous):

no i was still figuring it out

OpenStudy (anonymous):

@Jack1 i dont understand what to next in the equation i keep getting confused

OpenStudy (jack1):

ok, well lets back it up, and you tell me which step ur having trouble with and we'll go through it together so just say yes or no after each step, and well either keep going or explain more so 1. \(\Large V = \frac{L \times W \times H}3\) 2. "base of a square pyramid"... so from question, we know it's a square pyramid so the base is a square Length = Width new equation: \(\Large V = \frac{L \times L \times H}3\) all cool so far?

OpenStudy (anonymous):

got it (:

OpenStudy (jack1):

nice 3. from the question again: "Find the length .... if the volume is 576 cubic inches and has a height of 3 inches. " so find L if V = 576 and h = 3 put the values of V and H into our equation new equation: \(\Large V = \frac{L \times L \times H}3\) \(\Large 576 = \frac{L \times L \times 3}3\) still with me?

OpenStudy (anonymous):

sorry i was away from my computer but i understand still (:

OpenStudy (jack1):

all good 4. there's a 3 top and bottom in the fraction, so it's times 3 and then divide by 3... so they cancel eachother out \(\Large 576 = \frac{L \times L \times \color{red}{3}} {\color{red}{3}}\) \(\Large 576 = \frac{L \times L \times 1} 1\) \(\Large 576 = L \times L \) \(\Large 576 = L^2 \) all cool still?

OpenStudy (anonymous):

ohhh ok i understand thats were i was confused before (:

OpenStudy (jack1):

got it then, sweet! so are you ok to solve from here?

OpenStudy (anonymous):

yea i think so um.. for the last where it say 576=L/2 would then be 576= 2x2?

OpenStudy (jack1):

no, that's L squared, so 576 = L x L so if you take the square root of both sides you're left with : ____ (number?) = L

OpenStudy (anonymous):

oooh ok im sorry i feel like a dork now the 24=L correct (:

OpenStudy (jack1):

yep, perfect! ;)

OpenStudy (jack1):

and only dorks use the word dork :P

OpenStudy (anonymous):

lol thanks ^w^ you dont mind helping me with a few more questions do you ??

OpenStudy (jack1):

sure there's a lot of LaTex in this post tho, so maybe open a new one so it's faster?

OpenStudy (anonymous):

oh ok (: